Circular No. 7310

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999ex IN IC 5179 R. Martin, A. Williams, and S. Woodings, Perth Observatory, report

the discovery of a second supernova in IC 5179 (cf. IAUC 7272, SN 1999ee), found in an image taken on Nov. 9.51 UT in the course of the PARG Automated Supernova Search. Images obtained on Nov. 10.52 and 14 show that the object is slowly brightening. SN 1999ex is not visible in a deep exposure taken on Oct. 25.58 (estimated limiting mag 19). Position for SN 1999ex from three unfiltered CCD images on Nov. 10.53 with the Perth 0.25-m Mike Candy Telescope by J. Biggs and A. Verveer: α = 22^{h}16^{m}07^{s}.27 ± 0^{s}.02, δ = −36^{o}50′53′′.7 ± 0′′.3 (equinox 2000; mag ∼ 18.5 ± 1.0).

M. Hamuy, University of Arizona; and M. M. Phillips, Carnegie Observatories, report that a spectrum (range 350–1000 nm, resolution 0.3 nm) of SN 1999ex, taken on Nov. 14.13 UT with the New Technology Telescope at the European Southern Observatory (ESO), shows close resemblance to that of the type-Ic SN 1994I taken on 1994 Apr. 12 (3 days after V maximum; Filippenko et al. 1995, Ap.J. 450, L11). UBV images taken with the Cerro Tololo 0.91-m telescope by L. González, E. Rubenstein, P. Candia, and L. Strolger between Nov. 6 and 14 confirm the presence of SN 1999ex and reveal that maximum light occurred in all three bands on Nov. 13 at B = 17.3 and V = 16.6 (±0.1).

SUPERNOVA 1999ey IN ANONYMOUS GALAXY S. Benetti, Telescopio Nazionale Galileo, on behalf of the STRESS

team (E. Cappellaro, A. Pastorello, M. Prevedello, M. Salvo, and M. Turatto, Padova; J. Danziger, P. Mazzali and L. Rizzi, Trieste; and F. Patat, ESO), reports the discovery on a CCD frame taken on Nov. 10.17 UT of a supernova during a search performed with the ESO/MPI 2.2-m telescope (+ Wide Field Imager). SN 1999ey is located 0′′.5 west and 1′′.1 south of the nucleus of a faint galaxy, the latter being located at α = 0^{h}58^{m}03^{s}.39, δ = −27^{o}40′32′′.3 (equinox 2000.0). The new object was not present on Sept. 13 down to a limiting mag V = 23.5. SN 1999ey was confirmed with the ESO 3.6-m telescope (+ EFOSC2) on Nov. 14.20 at V ∼ 20.3, and a spectrum (range 350–1000 nm, resolution 2 nm) shows the features of a type-IIn supernova with strong H I Balmer emissions superimposed on a relatively blue continuum. Hα shows a composite profile, consisting of a narrow, barely resolved component (FWHM ∼ 600 km/s) on top of a broader one (FWHM ∼ 3350 km/s). The redshift, as measured from the narrow Hα emission of the supernova, is z ∼ 0.093.

1999 November 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green